# SimuCourt: Building Judicial Decision-Making Agents with Real-world Judgement Documents

### v1

@@ Appendix E Retrieval module
Result 1: The defendant shall return the principal amount of 800,000 yuan to the plaintiff within ten days from the effective date of this judgment, and pay interest based on the unpaid principal; Result 2: The defendant shall pay the plaintiff’s attorney fees of 49,000 yuan and the preservation guarantee fee of 1,800 yuan within ten days from the effective date of this judgment.
Table 10: Civil case example.
Cause of action Item Content
Labor and Social Security Administration
Case analysis
The court finds that the plaintiff, a company, and the third party, Wang, have a clear labor contract relationship with defined rights and obligations. The fact that Wang was injured in an accident during working hours and at the workplace due to work-related reasons is clear and well-evidenced. The local authority of Gangcheng District, upon receiving the company’s application for Wang’s work-related injury recognition and legally reviewing the relevant materials, made a decision on the work-related injury recognition within 60 days and delivered the decision document, in compliance with Articles 14(1), 19, and 20 of the Work Injury Insurance Regulations, with legal procedures followed.
Legal grounds
Article 69 of the Administrative Litigation Law of the People’s Republic of China; Paragraph 1 of Article 14 of the Work Injury Insurance Regulations; Article 19 of the Work Injury Insurance Regulations; Article 20 of the Work Injury Insurance Regulations.
Judgement
Result: Dismiss the plaintiff’s claim.
Table 11: Administrative case example.
Model Legal Grounds Judgement Results
Civil and Admini. Charge prison term Fine
SimuCourt 0.203 0.507 0.887 0.337 0.500
w/o Court simulation 0.171 0.473 0.875 0.300 0.462
w/o Knowledge base 0.145 0.462 0.850 0.312 0.475
w/o Web search 0.196 0.488 0.865 0.325 0.487
Table 12: Ablation study

### v3

@@ E.2 Judgement Evaluation for Civil and Administrative Cases
Result1: The defendant, Company XYZ, shall return the plaintiff’s rental deposit of 6000 yuan and the remaining rent of 17025 yuan within five days from the effective date of this judgment. Result1: The defendant shall return the deposit of 6000 yuan to the plaintiff.
Result2: The defendant, Company XYZ, shall pay the plaintiff a penalty of 5950 yuan within five days from the effective date of this judgment. Result2: The defendant shall return the remaining rent of 17025 yuan to the plaintiff.
Result3: The plaintiff’s other claims are dismissed
The analysis by GPT-4:
Key Point Match Evaluation The standard answer contains 4 key points: the return of the rental deposit of 6000 yuan, the return of the remaining rent of 17025 yuan, the payment of a penalty of 5950 yuan, and the dismissal of other claims. The generated answer’s "Result 1" matches the first key point (the return of the rental deposit of 6000 yuan), and "Result 2" matches the second key point (the return of the remaining rent of 17025 yuan). Therefore, the score for key point match is 2.
Answer Accuracy Evaluation The generated answer consists of 2 results. The "Result 1" of the generated answer accurately matches one key point without including incorrect information, thus scoring 1 point. Similarly, the "Result 2" of the generated answer accurately matches another key point without including incorrect information, also scoring 1 point. Therefore, the score for accuracy is 2.
Output Result Standard Key Points Count: 4, Key Point Match Score: 2, Generated Answer Results Count: 2, Answer Accuracy Score: 2
Precision Answer Accuracy Score / Generated Answer Results Count = 2/2
Recall Key Point Match Score / Standard Key Points Count = 2/4
Table 14: A prompt demonstration for GPT-4 evaluator (translated from Chinese).